OpenStudy (anonymous):

Alice has a basket containing 6 red socks, 4 blue socks, and 12 white socks. Find the theoretical probability that Alice will randomly select a red sock from the basket. Select the best answer from the choices provided. 6/16 1/2 3/11 1/22

OpenStudy (anonymous):

ok so u know liitle bit about probability right

OpenStudy (anonymous):

can u plz tell me how many total sockes we have?

OpenStudy (anonymous):

6+4+12=22 socks

OpenStudy (anonymous):

now we need to find what is probability we ll have a red socks which are only 6 in total of 22 socks

OpenStudy (anonymous):

so \[\frac{ 6 }{ 22 } =\frac{ 3 }{ 11 }\]

OpenStudy (anonymous):

Thank you @ayeshaafzal221